# A Hybrid Iterative Neural Solver Based on Spectral Analysis for Parametric PDEs

### v2

@@ Abstract
Recently, deep learning-based hybrid iterative methods (DL-HIM) have emerged as a promising approach for designing fast neural solvers to tackle large-scale sparse linear systems. DL-HIM combine the smoothing effect of simple iterative methods with the spectral bias of neural networks, which allows them to effectively eliminate both high-frequency and low-frequency error components. However, their efficiency may decrease if simple iterative methods can not provide effective smoothing, making it difficult for the neural network to learn mid-frequency and high-frequency components. This paper first conducts a convergence analysis for general DL-HIM from a spectral viewpoint, concluding that under reasonable assumptions, DL-HIM exhibit a convergence rate independent of grid size $h$ and physical parameters $\boldsymbol{\mu}$. To meet these assumptions, we design a neural network from an eigen perspective, focusing on learning the eigenvalues and eigenvectors corresponding to error components that simple iterative methods struggle to eliminate. Specifically, the eigenvalues are learned by a meta subnet, while the eigenvectors are approximated using Fourier modes with a transition matrix provided by another meta subnet. The resulting DL-HIM, termed the Fourier Neural Solver (FNS), can be trained to achieve a convergence rate independent of grid size and PDE parameters by designing a loss function that ensures the neural network complements the smoothing effect of Jacobi iterative methods. We verify the performance of FNS on six types of linear parametric PDEs.
@@ 1 Introduction
are ubiquitous in scientific and engineering applications, particularly those arising from the discretization of partial differential equations (PDEs). Developing efficient, robust, and adaptive numerical methods to solve these equations remains a significant challenge for researchers in applied mathematics. Iterative methods [1] are effective methods for solving such systems. Starting with an initial guess $\boldsymbol{u}^{(0)}$, the simplest form is relaxation
@@
The primary motivation is that $\mathcal{H}$ tends to fit low-frequency functions due to spectral bias [33, 34, 35], while simple preconditioners $\boldsymbol{B}$, such as damped Jacobi and Gauss-Seidel, are effective for eliminating high-frequency error components. We refer to (4a) as the smoothing iteration because it is primarily intended to eliminate oscillatory error components. However, it can not fully achieve this goal in certain problems, so the term “smoothing iteration” is used in a broader sense, as long as it can reduce some parts of the error. The neural network $\mathcal{H}:\mathbb{C}^{N}\rightarrow\mathbb{C}^{N}$ need be applicable for any $N\in\mathbb{Z}^{+}$. Therefore, it is typically designed as a discretization-invariant neural operator. The iterative method (4) was first proposed in [36], where DeepONet [37] was used as $\mathcal{H}$ and combined with different preconditioners $\boldsymbol{B}$ to test the compressibility with (4) for error components of various frequencies. This approach, referred to as the hybrid iterative numerical transferable solver (HINTS), was later extended to handle different geometries [38] and to solve the Helmholtz equation [39]. Around the same time, the authors of [40] employed local Fourier analysis (LFA) [41] to estimate which frequency components the selected $\boldsymbol{B}$ could effectively eliminate. Inspired by the fast Poisson solver [42, 43], they designed $\mathcal{H}$ based on the fast Fourier transform (FFT) to eliminate error components difficult for $\boldsymbol{B}$ to handle. In 2023, the authors of [44] designed a multilevel structure that mimics MG as $\mathcal{H}$, with shared parameters across different levels. This approach demonstrated good computational efficiency for convection-diffusion equations. In 2024, the authors of [45] used the SNO [46] as $\mathcal{H}$ and further accelerated it with FCG. In the same year, the authors of [47] used MIONet [48] as $\mathcal{H}$ and analyzed the convergence rate of the corresponding DL-HIM for the Poisson equation for the first time. Also in 2024, the author of [49] used a graph neural network to directly approximate $\boldsymbol{A}^{-1}$ as a nonlinear preconditioner for FGMRES.
@@
The main contents of this paper are as follows. First, we do a convergence analysis for the general DL-HIM (4) from a spectral viewpoint. For a diagonalizable matrix $\boldsymbol{A}$, under reasonable assumptions on $\boldsymbol{B}$ and $\mathcal{H}$, we can obtain a DL-HIM with a convergence rate independent of the mesh size $h$ and physical parameters $\boldsymbol{\mu}$. To ensure that $\mathcal{H}$ meets its assumptions, we then design $\mathcal{H}$ from an eigen perspective. Building on our previous work [40], we improve $\mathcal{H}$ by replacing the use of a Fourier matrix as the eigenvector matrix with a transition matrix that maps from the Fourier modes to the eigenvector basis. By introducing two meta subnets to learn the inverse of the eigenvalues and the transition matrix, we transform the requirements on $\mathcal{H}$ into requirements on the meta subnets. This approach avoids the constraints of spectral bias, enabling $\mathcal{H}$ to approximate not only low-frequency error components but also error components that $\boldsymbol{B}$ struggles to eliminate. The improved solver, still called Fourier Neural Solver (FNS), is implemented as an end-to-end, differentiable neural solver for linear systems derived from structured grids. Finally, we design a reasonable loss function and construct easily accessible and targeted training data to train FNS and test its performance on a variety of classical PDE discrete systems. In the numerical experiments, we first verify the validity of $\mathcal{H}$ for the Poisson equation using analytic eigenvalues and eigenvectors. This results in an FNS where the convergence rate remains independent of the problem size. For random diffusion equations, the selected $\boldsymbol{B}$ effectively eliminates high-frequency errors, while a well-trained $\mathcal{H}$ can learn low-frequency and mid-frequency error components. This enables FNS to achieve a convergence rate independent of both physical parameters and grid size. For multi-scale problems, such as discrete systems arising from anisotropic diffusion, strong convection, and jumping diffusion equations, FNS maintains a convergence rate independent of physical parameters and grid size on small and medium scales with additional support. In the case of the Helmholtz equation, even though the chosen $\boldsymbol{B}$ amplifies the lowest-frequency errors, the trained $\mathcal{H}$ effectively corrects these errors, ensuring a convergent FNS. In summary, the advantages of FNS over other DL-HIM are as follows:
@@
    FNS can handle error components with different frequencies. Several DL-HIMs [36, 38, 47, 39, 45] train $\mathcal{H}$ independently, following an approach similar to that used in neural operators. However, the trained $\mathcal{H}$ tends to capture only low-frequency components due to spectral bias. In contrast, the $\mathcal{H}$ of FNS, trained in Fourier space, is not affected by spectral bias. It can also automatically learn error components that are difficult for $\boldsymbol{B}$ to eliminate because it is trained within (4) in an end-to-end manner.
@@
The rest of this paper is organized as follows: Section 2 establishes a convergence analysis framework for the general DL-HIM; Section 3 proposes FNS and designs reasonable training data and loss function under the guidance of convergence analysis; Section 4 conducts numerical experiments on four types of second-order linear PDEs to verify the theoretical analysis; and Section 5 gives a summary and outlook.
@@ 3 Fourier neural solver
Remark 2
In fact, we did not extensively explore the architecture of the meta subnets. Hyper nets in [22, 40], encoder nets in [57, 44], and branch nets in [36, 47] play similar roles but utilize different network structures. By tailoring the meta structure to the specific characteristics of the problem, one might achieve better results than those presented in the next section.
@@ 4 Numerical experiments
    Poisson equation: Symmetric, positive definite.
@@
    Random diffusion equation: Poisson-like elliptic problem, non-symmetric, positive definite.
@@
    Anisotropic diffusion equation: Multiscale, symmetric positive definite.
@@
    Jumping diffusion equation: Multiscale with different scales at different grid points, non-symmetric, positive definite.
@@ 4.2 Random diffusion equations
Next, we train FNS to ensure $\mathcal{H}$ satisfies assumption (30) as much as possible. For this problem, we aim for the trained FNS to generalize well for both $a$ and $f$. According to previous analysis, the asymptotic convergence rate of FNS is independent of RHS (as we will verify below), so $a$ is the only parameter of interest, corresponding to $\boldsymbol{\mu}$ in PPDE (5). We generate 10,000 random diffusion coefficients $a_{i}$ with scale $n=63$ and sample a random RHS $\boldsymbol{f_{i}}\sim\mathcal{N}(0,\boldsymbol{I})$ for each parameter $a_{i}$. We employ the loss function (31) for unsupervised training, utilize the Adam optimizer with an initial learning rate of $10^{-4}$ and a learning rate schedule that halves the learning rate every 100 epochs. Additionally, we increase $K$ by one every 100 epochs and set $M=10$. The training loss is depicted in Figure 8.
@@
Then we test the convergence speed of FNS on different scales. For each scale, we sample 10 different coefficients $a_{i}$ and fix $f=1$. Figure 10 illustrates the mean and standard deviation of the FNS iteration counts required to achieve a relative residual of less than $10^{-6}$. The results indicate that when $n\leq 255$, the convergence rate of FNS is almost independent of the grid size $h$ and the coefficients $a$, which aligns with our analysis. However, as the scale increases further, the number of required iterations grows linearly with the scale.
Figure 10: FNS iteration counts to satisfy $\|\boldsymbol{r}^{(k)}\|/\|\boldsymbol{f}\|<10^{-6}$, starting with zero initial value, for different random diffusion equations. (a) Trained on $n=63$ but solve problems with different scales; (b) Trained on different grid sizes but test on $n=511$.
To address this issue, we keep the parameter counts of the meta subnets invariant while increasing the training data. According to the neural scaling law, as the size of the training data increases, the performance of FNS improves accordingly. Figure 10 presents the FNS iteration counts trained on different grid sizes but tested on a scale of $n=511$. As the training data size increases, FNS performance improves, indicating that a fully trained FNS better satisfies (30), making the number of iterations independent of the scale. However, training with large datasets is computationally expensive, so enhancing the generalization capability of FNS trained on smaller scales remains a challenge to address in the near future.
@@ 4.3 Anisotropic diffusion equations
Similarly, we test the required iteration counts of FNS when solving different anisotropic equations with various parameters $\boldsymbol{\mu}$ on different mesh scales $h$. For each scale, we sample 10 different $\boldsymbol{\mu}$ and calculate the mean and standard deviation of iteration counts. Since smaller $\xi$ leads to a stronger multi-scale property of the discrete system, we sample the test parameters within $[10^{-7},10^{-5}]\times[-\pi,\pi]$. Figure 14 shows the results. It can be seen that when $n\leq 127$, the FNS iteration counts are almost invariant with $h$ and $\boldsymbol{\mu}$, but they increase significantly beyond $n=127$. We believe this may be addressed by increasing the amount of large-scale training data, increasing the parameters of meta subnets, and using more powerful optimization algorithms.
Figure 14: FNS iteration counts to satisfy $\|\boldsymbol{r}^{(k)}\|/\|\boldsymbol{f}\|<10^{-6}$, starting with zero initial value, for anisotropic diffusion equations with different scales.
We now compare the performance of FNS with HINTS, which employs DeepONet as $\mathcal{H}$. For random diffusion equations (32), the damped Jacobi smoother $\boldsymbol{B}$ effectively eliminates high-frequency errors, while DeepONet efficiently learns low-frequency errors, leveraging the frequency principle. Consequently, HINTS can be trained to achieve fast convergence on small scales, as shown in Figure 15(a). However, for anisotropic diffusion problems, where the smoother $\boldsymbol{B}$ struggles to eliminate high-frequency errors along the anisotropic direction, DeepONet also encounters difficulties in learning these errors due to the frequency principle. Numerical experiments show that HINTS converges slowly under these circumstances, as illustrated in Figure 15(b). These findings validate the assertion that existing DL-HIMs face challenges in effectively correcting errors that $\boldsymbol{B}$ fails to eliminate.
@@
Figure 15: Convergence histories of FNS and HINTS for solving random and anisotropic diffusion equations, $n=63$.
@@ 4.5 Jumping diffusion equations
Figure 18(a)18(b) presents an example of the coefficients and the corresponding solution, where the red region represents $\Omega_{1}$ and the blue region represents $\Omega_{2}$.
@@ 4.6 Helmholtz equation
$-\Delta u-k^{2}u&=f,\text{ in }\Omega=(0,1)2,\\ \nabla u\cdot\mathbf{n}-iku&=0,\text{ on }\partial\Omega,\\ \end{aligned}\right.$ (42)
@@ 5 Conclusion and discussion
FNS provides a versatile framework for solving general linear systems. However, achieving theoretical results for complex problems requires leveraging the problem characteristics and developing more suitable meta network architectures along with more efficient training algorithms. In the near future, we aim to enhance our approach to better solve multi-scale and indefinite problems. Additionally, we plan to extend the application of DL-HIM to unstructured grids and 3D problems. Furthermore, we intend to build a comprehensive software library on DL-HIM to integrate all related methods within the same framework, facilitating easier implementation and comparison. Through continuous efforts, we hope to make DL-HIM a widely used preconditioning method similar to MG and DDM, and a neural pde solver as popular as PINN and neural operators.

### v3

@@ Abstract
Deep learning-based hybrid iterative methods (DL-HIM) have emerged as a promising approach for designing fast neural solvers to tackle large-scale sparse linear systems. DL-HIM combine the smoothing effect of simple iterative methods with the spectral bias of neural networks, which allows them to effectively eliminate both high-frequency and low-frequency error components. However, their efficiency may decrease if simple iterative methods can not provide effective smoothing, making it difficult for the neural network to learn mid-frequency and high-frequency components. This paper first conducts a convergence analysis for general DL-HIM from a spectral viewpoint, concluding that under reasonable assumptions, DL-HIM exhibit a convergence rate independent of grid size $h$ and physical parameters $\boldsymbol{\mu}$. To meet these assumptions, we design a neural network from an eigen perspective, focusing on learning the eigenvalues and eigenvectors corresponding to error components that simple iterative methods struggle to eliminate. Specifically, the eigenvalues are learned by a meta subnet, while the eigenvectors are approximated using Fourier modes with a transition matrix provided by another meta subnet. The resulting DL-HIM, termed the Fourier Neural Solver (FNS), can be trained to achieve a convergence rate independent of PDE parameters and grid size within a local neighborhood of the training scale by designing a loss function that ensures the neural network complements the smoothing effect of the damped Jacobi iterative methods. We verify the performance of FNS on five types of linear parametric PDEs.
@@ 1 Introduction
are ubiquitous in scientific and engineering applications, particularly those arising from the discretization of partial differential equations (PDEs). Developing efficient, robust, and scalable numerical methods to solve these equations remains a significant challenge for researchers in applied mathematics. Iterative methods [1] are effective methods for solving such systems. Starting with an initial guess $\boldsymbol{u}^{(0)}$, the simplest form is relaxation
@@
The primary motivation is that $\mathcal{H}$ tends to fit low-frequency functions due to spectral bias [33, 34, 35], while simple preconditioners $\boldsymbol{B}$, such as damped Jacobi and Gauss-Seidel, are effective for eliminating high-frequency error components. We refer to (4a) as the smoothing iteration because it is primarily intended to eliminate oscillatory error components. However, it can not fully achieve this goal in certain problems, so the term “smoothing iteration” is used in a broader sense, as long as it can reduce some parts of the error. The neural network $\mathcal{H}:\mathbb{C}^{N}\rightarrow\mathbb{C}^{N}$ need be applicable for any $N\in\mathbb{Z}^{+}$. Therefore, it is typically designed as a discretization-invariant neural operator. The iterative method (4) was first proposed in [36], where DeepONet [37] was used as $\mathcal{H}$ and combined with different preconditioners $\boldsymbol{B}$ to test the compressibility for error components of various frequencies. This approach, referred to as the hybrid iterative numerical transferable solver (HINTS), was later extended to handle different geometries [38] and to solve the Helmholtz equation [39]. Around the same time, the authors of [40] employed local Fourier analysis (LFA) [41] to estimate which frequency components the selected $\boldsymbol{B}$ could effectively eliminate. Inspired by the fast Poisson solver [42, 43], they designed $\mathcal{H}$ based on the fast Fourier transform (FFT) to eliminate error components difficult for $\boldsymbol{B}$ to handle. In 2023, the authors of [44] designed a multilevel structure that mimics MG as $\mathcal{H}$, with shared parameters across different levels. This approach demonstrated good computational efficiency for convection-diffusion equations. In 2024, the authors of [45] used the SNO [46] as $\mathcal{H}$ and further accelerated it with FCG. In the same year, the authors of [47] used MIONet [48] as $\mathcal{H}$ and analyzed the convergence rate of the corresponding DL-HIM for the Poisson equation for the first time. Also in 2024, the author of [49] used a graph neural network to directly approximate $\boldsymbol{A}^{-1}$ as a nonlinear preconditioner for FGMRES.
@@
The main contents of this paper are as follows. First, we do a convergence analysis for the general DL-HIM (4) from a spectral viewpoint. For a diagonalizable matrix $\boldsymbol{A}$, under reasonable assumptions on $\boldsymbol{B}$ and $\mathcal{H}$, we can obtain a DL-HIM with a convergence rate independent of the mesh size $h$ and physical parameters $\boldsymbol{\mu}$. To ensure that $\mathcal{H}$ meets its assumptions, we then design $\mathcal{H}$ from an eigen perspective. Building on our previous work [40], we improve $\mathcal{H}$ by replacing the use of a Fourier matrix as the eigenvector matrix with a transition matrix that maps from the Fourier modes to the eigenvector basis. By introducing two meta subnets to learn the inverse of the eigenvalues and the transition matrix, we transform the requirements on $\mathcal{H}$ into requirements on the meta subnets. This approach avoids the constraints of spectral bias, enabling $\mathcal{H}$ to approximate not only low-frequency error components but also error components that $\boldsymbol{B}$ struggles to eliminate. The improved solver, still called Fourier Neural Solver (FNS), is implemented as an end-to-end, differentiable neural solver for linear systems derived from structured grids. Finally, we design a reasonable loss function and construct easily accessible and targeted training data to train FNS and test its performance on a variety of classical PDE discrete systems. In the numerical experiments, we first verify the validity of $\mathcal{H}$ for the Poisson equation using analytic eigenvalues and eigenvectors. This results in an FNS where the convergence rate remains independent of the problem size. For random diffusion equations, the selected $\boldsymbol{B}$ effectively eliminates high-frequency errors, while a well-trained $\mathcal{H}$ can learn low-frequency and mid-frequency error components. This enables FNS to achieve a convergence rate independent of both physical parameters and grid size within a local neighborhood of the training scale. For multi-scale problems, such as discrete systems arising from anisotropic diffusion, convection-diffusion, and jumping diffusion equations, FNS maintains a convergence rate independent of physical parameters and grid size on medium scales. In the case of the Helmholtz equation, even though the chosen $\boldsymbol{B}$ amplifies the lowest-frequency errors, the trained $\mathcal{H}$ effectively corrects these errors, ensuring a convergent FNS. In summary, the advantages of FNS over other DL-HIM are as follows:
@@
    FNS can handle error components with different frequencies. Several DL-HIMs [36, 38, 47, 39, 45] train $\mathcal{H}$ independently, following an approach similar to that used in neural operators. However, the trained $\mathcal{H}$ tends to capture only low-frequency components due to spectral bias. In contrast, FNS mitigates this issue by learning in the frequency domain and employing an end-to-end training strategy, which enables better performance.
@@
The rest of this paper is organized as follows: Section 2 establishes a convergence analysis framework for the general DL-HIM; Section 3 proposes FNS and designs reasonable training data and loss function under the guidance of convergence analysis; Section 4 conducts numerical experiments on several types of second-order linear PDEs to verify the theoretical analysis; and Section 5 gives a summary and outlook.
@@ 4 Numerical experiments
    Poisson equation: SPD.
@@
    Random diffusion equation: Poisson-like elliptic problem, SPD.
@@
    Anisotropic diffusion equation: Multiscale, SPD.
@@
    Jumping diffusion equation: Multiscale with different scales at different grid points, SPD.
@@ 4.2 Random diffusion equations
Next, we train FNS to ensure $\mathcal{H}$ satisfies assumption (30) as much as possible. For this problem, we aim for the trained FNS to generalize well for both $a$ and $f$. According to previous analysis, the asymptotic convergence rate of FNS is independent of RHS (as we will verify below), so $a$ is the only parameter of interest, corresponding to $\boldsymbol{\mu}$ in PPDE (5). We generate 10,000 random diffusion coefficients $a_{i}$ and sample a random RHS $\boldsymbol{f_{i}}\sim\mathcal{N}(0,\boldsymbol{I})$ for each parameter $a_{i}$. We employ the loss function (31) for unsupervised training, utilize the Adam optimizer with an initial learning rate of $10^{-4}$ and a learning rate schedule that halves the learning rate every 100 epochs. Additionally, we increase $K$ by one every 100 epochs. The training loss is depicted in Figure 8.
@@
We then evaluated the performance of FNS across different discretization scales. For testing, we sampled 10 distinct coefficients $a_{i}$ and fixed the RHS $f=1$. Table 1 presents the mean and standard deviation of the iteration counts required by FNS (trained at specific scales) to reduce the relative residual below $10^{-6}$. The results reveal the following:
  1. 1.
    FNS achieves optimal performance at the training scale, with iteration counts increasing at both smaller and larger scales.
  2. 2.
    FNS exhibits parameter-independent convergence that is only weakly dependent on discretization scale within a certain range near the training scale.
  3. 3.
    FNS generalizes better to scales smaller than the training scale than to significantly larger scales.
These observations confirm that while FNS achieves scale-independent convergence near the training regime, it does not yet demonstrate universal scale invariance. At present, this issue can be mitigated by training on larger-scale problems. Developing more cost-effective approaches to enhance scale generalization will be a focus of our future work.
Table 1: FNS iteration counts for FEM discretizations of random diffusion equations at different grid size. Each entry represents the mean $\pm$ std over 10 random coefficient samples, with a fixed RHS $f=1$.
Grid size $n$ 31 63 127 255 511
Model trained at $n=63$ $15.6\pm 1.74$ $\mathbf{12.3\pm 0.78}$ $13.7\pm 0.78$ $25.0\pm 1.34$ $45.2\pm 3.67$
Model trained at $n=255$ $15.0\pm 2.88$ $14.3\pm 2.05$ $12.6\pm 1.43$ $\mathbf{12.5\pm 1.43}$ $15.9\pm 1.14$
@@ 4.3 Anisotropic diffusion equations
We next evaluate the performance of FNS in solving anisotropic diffusion equations with varying parameters $\boldsymbol{\mu}$ across different mesh resolutions $h$. FNS was trained separately at $63^{2}$ and $511^{2}$ grids. Due to the high training cost at $511^{2}$, the training dataset for this case was limited to 2,000 samples (200 $\boldsymbol{\mu}_{i}$, each paired with 10 different RHS $\boldsymbol{f}_{i}$). Table 2 reports the iteration counts of FNS trained at different resolutions. It can be observed that:
  1. 1.
    FNS trained at $n=63$ exhibits mesh- and parameter-independent convergence for test grids with $n\in[31,127]$, but its performance deteriorates on larger grids ($n\geq 255$).
  2. 2.
    FNS trained at $n=511$ achieves comparable convergence performance at its training scale to that of the $n=63$ model at its own scale.
  3. 3.
    The $n=511$-trained FNS performs worse than the $n=63$ model on smaller grids, which may be attributed to limited training data at large scales.
Table 2: FNS iteration counts for anisotropic diffusion equations across grid sizes. Each entry reports the mean $\pm$ standard deviation over 10 random coefficients with $\log_{10}\varepsilon\sim\mathcal{U}(-6,0)$, $\theta=\frac{5\pi}{12}$, and $f=1$.
Grid size $n$ 31 63 127 255 511
Model trained at $n=63$ $22.7\pm 2.45$ $\mathbf{21.5\pm 1.97}$ $24.4\pm 2.96$ $78.5\pm 10.86$ $>200$
Model trained at $n=511$ $78.3\pm 9.53$ $99.5\pm 8.02$ $83.6\pm 5.16$ $78.0\pm 10.69$ $\mathbf{24.5\pm 6.44}$
We then compare the performance of FNS with HINTS, which employs DeepONet as $\mathcal{H}$. For the random diffusion equation (32), the damped Jacobi smoother $\boldsymbol{B}$ effectively removes high-frequency errors, while DeepONet efficiently learns the remaining low-frequency components due to the spectral bias. As a result, HINTS achieves rapid convergence on small-scale problems, as shown in Figure 13(a). However, for anisotropic diffusion equations, where $\boldsymbol{B}$ struggles to eliminate high-frequency errors aligned with the anisotropy direction, DeepONet also fails to learn such errors effectively, again due to the spectral bias. In this case, HINTS converges slowly, as illustrated in Figure 13(b). In contrast, FNS effectively mitigates spectral bias by operating in the frequency domain and adopting an end-to-end training strategy. This design enables the network to better capture a broader range of error components, including those that are typically challenging for conventional neural solvers, thereby achieving more robust and efficient convergence.
@@
Figure 13: Convergence histories of FNS and HINTS for solving random and anisotropic diffusion equations, with $n=63$.
@@ 4.5 Jumping diffusion equations
Figure 16(a)16(b) presents an example of the coefficients and the corresponding solution, where the red region represents $\Omega_{1}$ and the blue region represents $\Omega_{2}$.
@@ 4.6 Helmholtz equation
$-\Delta u-k^{2}u&=f,\text{ in }\Omega=(0,1)^{2},\\ \nabla u\cdot\mathbf{n}-iku&=0,\text{ on }\partial\Omega,\\ \end{aligned}\right.$ (42)
@@ 5.1 Conclusions
5.2 Scope, limitations, and future work
On the scale-independence of FNS convergence rates. Numerical experiments on random and anisotropic diffusion equations indicate that FNS achieves convergence rates that are largely independent of discretization scale when tested on systems that are either smaller than or moderately larger than (e.g., up to twice the size of) the training scale. However, when the test scale significantly exceeds the training scale, the number of iterations required increases substantially. These findings suggest that FNS exhibits scale-independent convergence primarily within a local neighborhood of the training scale, but has not yet achieved universal scale invariance. Incorporating training data from larger-scale problems can improve generalization and enable convergence performance at extended scales comparable to that at smaller ones. On the other hand, this limitation may also be influenced by the choice of Meta-network architecture. We experimented with both UNet and FNO as Meta-$\lambda$. While FNO generally outperformed UNet and is therefore adopted in this work, UNet achieved better results in the case of random diffusion equations. This suggests that the Meta-network architecture plays a non-negligible role in determining convergence behavior. Designing more effective Meta-network architectures to improve scalability and generalization of FNS remains a key direction for future research.
On the consumption of training resources. The experiments in this paper are mainly conducted at the grid scale of $n=63$, which is a representative training scale based on a balance between computational efficiency and model performance. Table 5 summarizes the training resource consumption observed on an NVIDIA A100 GPU for various problem sizes, providing justification for this choice:
  1. 1.
    Hardware utilization: At $n=63$,
    • Computational resources are effectively utilized. Increasing $n$ from 31 to 63 leads to a VRAM usage increase of less than 3x and a per-epoch training time increase of only 2x, indicating underutilization at smaller scales.
    • Further increasing $n$ beyond 63 results in a more than 3x increase in training time, suggesting that $n=63$ achieves near-optimal parallelization efficiency on our hardware.
  2. 2.
    Acceptable training cost: Training at $n=63$ remains computationally feasible while producing performant models.
Table 5: Training resource consumption on an NVIDIA A100 GPU for different problem sizes of random diffusion equations. Each configuration uses $K=3$ in the loss function.
Grid size $n$ Data size Batch size VRAM usage (MB) Time per epoch (s)
31 $10^{5}$ 40 5837 21
63 $10^{5}$ 40 14179 40
127 $10^{5}$ 40 44083 125
Since the current network architecture indeed demonstrates strong generalization capability primarily near the training scale. To address large-scale computational demands, scaling up the training regime—i.e., training on larger problem sizes—becomes imperative. This is a key direction in our ongoing work, aimed at improving the robustness and scalability of FNS across a broader range of discretization levels.
On the applicability of FNS to nonsymmetric PDEs. The design of the operator $\mathcal{H}$ is motivated by the eigendecomposition of $\boldsymbol{A}$, which fundamentally assumes that $\boldsymbol{A}$ is symmetric or normal. Nevertheless, our numerical experiments in Section 4.4 on convection-diffusion equations demonstrate that FNS retains weak dependence on the discretization scale and achieves near-linear runtime, even when applied to nonsymmetric SUPG discretizations. Extending this capability to more general nonsymmetric discrete systems remains an important direction for future research.
On the applicability of FNS to general discrete systems. The discrete systems considered in this work are all derived from structured grids on regular domains. The applicability of FNS to more general cases—such as discrete systems arising from irregular domains, unstructured grids, three-dimensional PDEs, or PDE systems—remains an open question. A key challenge is the current reliance of FNS on the FFT, which assumes a regular and consistent grid structure. Extending the framework to accommodate arbitrary discretizations will be an important direction for future work, potentially through the integration of graph neural networks or geometry-aware architectures.